There is some money with Arun and some money with Kiran. If Arun gives Rs 60 to Kiran, then the amounts with them would be equal. If instead Kiran gives Rs 20 to Arun, then Arun would have Rs 160 less than twice the amount with Kiran. Find the amount of Kiran.
Answer by ikleyn(52786)

.
The given part translates to these two equations

A - 60 = K + 60,             (1)
2*(K - 20) = A + 20 + 160    (2)


Simplify and write in standard form

A = K + 120              
2K - 40 = A + 180,

or

A = K + 120                   (3)            
2K = A + 220,                 (4)


Next substitute (3) into (4) to get

2K = (K+120) + 220  ====>  K = 340.


Answer.  Kiran amount is Rs 340.


Check.   Then Arun's amount is A = K + 120 = 340 + 120 = 460.


                Then (1) takes the form  460-60 = 400 = 340+60        ! Correct !

                 and (2) takes the form  2*(340-20) = 640 = 460+180   ! Correct !                 

Solved.

Very standard problem. To see other similar solved problems, look into the lesson
    - Virtual exchange between two persons
in this site.